For a standard normal distribution, find: P(0.61 < z < 2.92)
Answer #1

Saletion: P (6.612 222-92) = P (242.92) - P(Z20.61) 0.9982 -0.7291 P 0-61< Z4 2.92) = 0.2691 fo Use normal distribution table